Find the solution to this equation: 3x + y = 6; y + 2 = x
Answer by mananth(16946)

3x + y = 6; y + 2 = x
substitute x
3(y+2)+y=6
3y+6+y=6
4y=0
y=0
plug y
y+2=x
0+2=x
x=2